# A multi-messenger hierarchical triple merger gravitational-wave event pair GW190514-GW190521 inside AGN J124942.3 + 344929

### v1

@@ II.6 Models including primary masses
Following the considerations from previous literature, we take the primary masses of gravitational wave events into account in our hypothetical models [5]. The equations (12) and (13) become
@@ III Results
Figure 2: Probability distribution of the associated mass. The empty histograms, each normalized, illustrate the associated mass distribution of GW190514-GW190521 for the association model, both without (red) and with (green) the inclusion of the primary mass $m_{1}$. In contrast, the filled histograms represent the final mass distribution of GW190514 (blue) and the second mass distribution of GW190521 (yellow), as obtained from the GWTC-2.1. The gray line indicates the mass prior conditioned for the association model, which corresponds to the final mass distribution of the first generation binary black hole mergers in the AGN formation channel.
@@
Figure 3: Probability distribution of the primary mass. The empty histograms, each normalized, represent the primary mass distribution for the association model (green). For comparison, the filled histograms depict the primary mass distribution from the GWTC-2.1. The gray line indicates the mass prior conditioned for the association model. On the left, the distributions for GW190514 are displayed, with the prior conditioned corresponding to the primary mass distribution of first generation binary black hole mergers in the AGN formation channel. On the right, the distributions for GW190521 are shown, with the prior conditioned reflecting the primary mass distribution of hierarchical binary black hole mergers in the AGN formation channel.
Furthermore, we find that the primary masses of these two gravitational wave events are also crucial for enhancing the probability of the association compared to the coincidence [5]. This is because binary black hole mergers in AGN disks are generally expected to have more massive components compared to other channels, such as isolated binary evolution or star clusters, due to accretion and frequent hierarchical mergers [9, 10, 11]. Therefore, we also consider the primary masses ($m_{1}$) of the gravitational wave events to assess the multi-messenger coincidence significance for the three transients (see Section II.6). Figure 3 presents the posterior probability densities of the primary masses of the two gravitational wave events, comparing the distributions from the GWTC-2.1 [2] with those from the association model (labeled as ‘Association with $m_{1}$’, green). As shown in this figure, the association model with $m_{1}$ provides better constraints, slightly favoring a more massive black hole for GW190521 and a less massive one for GW190514, compared to the GWTC-2.1 distributions [2]. Notably, GW190521’s primary black hole mass falls within the mass gap with nearly 100% probability, strongly suggesting that it is a hierarchical merger [47], potentially a second (or higher) generation merger resulting from GW190514 in an AGN environment. Incorporating the primary mass information significantly strengthens the association hypothesis, increasing the Bayes factor by approximately two orders of magnitude, corresponding to a log Bayes factor of ${\rm log}\mathcal{B}=21.5$ and an odds ratio of $\mathcal{O}\sim$$12\times 10^{4}:$ 1. This result aligns with the expectation that more massive binary black hole mergers are more likely to occur in the AGN formation channel [9, 10, 11]. Additionally, we find that the inclusion of primary mass data allows slightly for better constraints on the associated mass between GW190514 and GW190521 (labeled as ‘Association with $m_{1}$’, green, in Fig. 2).
@@
Finally, we note that in addition to the three detection associations, there may also be scenarios where only two detections share a common source. For comparison, we also investigate the scenarios, which has three different combinations, corresponding to the hypotheses $\mathcal{H}_{2}$, $\mathcal{H}_{3}$, and $\mathcal{H}_{4}$ in Table 1. Table 1 presents the Bayes factors and odds for various association combinations of GW190514, GW190521, and ZTF19abanrhr, indicating that a common origin shared by the three transients is strongly favored based on the observed data sets, in contrast to the associations involving only two of them. In particular, the odds of $\mathcal{H}_{1}$ is $\sim$14 (3902) times greater than the odds of $\mathcal{H}_{2}$, $\mathcal{H}_{3}$, and $\mathcal{H}_{4}$ (considering the primary masses). The analysis results, which account for the localizations and masses of the three transients, provide significantly robust support for the hypothesis of a common origin ($\mathcal{H}_{1}$). Moreover, we reanalyze the potential association by including the spins of the gravitational wave events. However, we find that the spins do not contribute significant information, yielding a Bayes factor of approximately 1.03 for the spin-only model.
@@ IV Conclusions and discussion
In this work, we investigate the potential association between GW190514 and GW190521 as a hierarchical triple merger associated with ZTF19abanrhr, taking into account of sky position, distance, and mass of the sources using a Bayesian criterion. Although the previous study has shown that GW190514 and GW190521 are a plausible hierarchical triple merger, their analysis is not significant with $p$-value of $\sim$0.14 [3] (generally, $p$-value of $\leq$0.05 is considered ‘significant’). Here, we introduce a new approach and propose using multi-messenger [4] signals to analyze the probability of the GW190514-GW190521 hierarchical triple merger. Our analysis (considering the primary masses) results greatly increase the probability, finding that a log Bayes factor of ${\rm log}\mathcal{B}=16.8~{}(21.5)$, corresponding to an odds ratio of $\mathcal{O}\sim$$1199~{}(2\times 10^{4}):$ 1, assuming an astrophysical prior odds of $\mathcal{P}\sim 10^{-5}$. This suggest strong evidence for the first hierarchical triple merger associated with an electromagnetic counterpart in the AGN formation channel. The finding of the multi-messenger hierarchical triple merger in the AGN is crucial for understanding the formation mechanisms of massive black holes, the role of AGNs in hierarchical mergers, and the implications of multi-messenger astronomy
@@
We show that the GW190514-GW190521 pair, which is separated by only one week, further supports the scenario of a triple hierarchical merger chain if associated with ZTF19abanrhr: i) Such a short time frame between the two mergers presents significant challenges for the reformation of a binary system and its subsequent merger. However, AGNs provide an environment conducive to overcoming these challenges [48, 49]. Specifically, the number density of black holes in the migration traps [35] of AGN disks is notably high due to migration [50]. This high density facilitates the rapid reformation of a binary system from GW190514’s remnant black holes, enabling a swift second merger with the aid of the dense gas in AGN disks; ii) GW190521 may not have had sufficient time to circularize in such a short interval, suggesting that it is likely to be eccentric [51, 52]. This inference is supported by previous analyses indicating a non-zero eccentricity for GW190521 [3], and AGN disk environments are known to promote an excess of eccentric mergers [16]. Our current methodological framework explicitly excludes time delay and eccentricity considerations in the statistical analysis, as these require dedicated waveform development that will be addressed in subsequent studies. Furthermore, we anticipate environmental effects to induce luminosity distance corrections of $\lesssim 1\%$ — a systematic uncertainty. This is tiny compared to the uncertainty in intrinsic luminosity distance for each binary black hole observation (typically ±$\pm 30\%$). The impact warrants future investigation through targeted parameter estimation studies.
@@ V Acknowledgments
This work is supported by National Key R&\&&D Program of China (2020YFC2201400), the National Natural Science Foundation of China (grant No. 11922303), and the Fundamental Research Funds for the Central Universities, Wuhan University (grant No. 2042022kf1182). We use the gravitational wave datasets released with GWTC-2.1 [2], labeled as “cosm” and “IMRPhenomXPHM.” This research has made use of data or software obtained from the Gravitational Wave Open Science Center (https://gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan. This analysis was made possible following software packages: NumPy [53], SciPy [54], Matplotlib [55], emcee [56], IPython [57], corner [58], seaborn [59], and Astropy [60].

### v2

@@ II.6 Models including primary masses
Following the considerations from previous literature, we take the primary masses of GW events into account in our hypothetical models [5]. The equations (12) and (13) become
@@ III Results
Figure 2: Probability distribution of the associated mass. The empty histograms, each normalized, illustrate the associated mass distribution of GW190514-GW190521 for the association model, both without (red) and with (green) the inclusion of the primary mass $m_{1}$. In contrast, the filled histograms represent the final mass distribution of GW190514 (blue) and the second mass distribution of GW190521 (yellow), as obtained from the GWTC-2.1. The gray line indicates the mass prior conditioned for the association model, which corresponds to the final mass distribution of the first generation BBH mergers in the AGN formation channel.
@@
Figure 3: Probability distribution of the primary mass. The empty histograms, each normalized, represent the primary mass distribution for the association model (green). For comparison, the filled histograms depict the primary mass distribution from the GWTC-2.1. The gray line indicates the mass prior conditioned for the association model. On the left, the distributions for GW190514 are displayed, with the prior conditioned corresponding to the primary mass distribution of first generation BBH mergers in the AGN formation channel. On the right, the distributions for GW190521 are shown, with the prior conditioned reflecting the primary mass distribution of hierarchical BBH mergers in the AGN formation channel.
Furthermore, we find that the primary masses of these two GW events are also crucial for enhancing the probability of the association compared to the coincidence [5]. This is because BBH mergers in AGN disks are generally expected to have more massive components compared to other channels, such as isolated binary evolution or star clusters, due to accretion and frequent hierarchical mergers [9, 10, 11]. Therefore, we also consider the primary masses ($m_{1}$) of the GW events to assess the multi-messenger coincidence significance for the three transients (see Section II.6). Figure 3 presents the posterior probability densities of the primary masses of the two GW events, comparing the distributions from the GWTC-2.1 [2] with those from the association model (labeled as ‘Association with $m_{1}$’, green). As shown in this figure, the association model with $m_{1}$ provides better constraints, slightly favoring a more massive BH for GW190521 and a less massive one for GW190514, compared to the GWTC-2.1 distributions [2]. Notably, GW190521’s primary BH mass falls within the mass gap with nearly 100% probability, strongly suggesting that it is a hierarchical merger [48, 49, 50, 51, 52, 53, 54, 55], potentially a second (or higher) generation merger resulting from GW190514 in an AGN environment. Incorporating the primary mass information significantly strengthens the association hypothesis, increasing the Bayes factor by approximately two orders of magnitude, corresponding to a log Bayes factor of ${\rm log}\mathcal{B}=21.5$ and an odds ratio of $\mathcal{O}\sim$$12\times 10^{4}:$ 1. This result aligns with the expectation that more massive BBH mergers are more likely to occur in the AGN formation channel [9, 10, 11]. Additionally, we find that the inclusion of primary mass data allows slightly for better constraints on the associated mass between GW190514 and GW190521 (labeled as ‘Association with $m_{1}$’, green, in Fig. 2).
@@
Finally, we note that in addition to the three detection associations, there may also be scenarios where only two detections share a common source. For comparison, we also investigate the scenarios, which has three different combinations, corresponding to the hypotheses $\mathcal{H}_{2}$, $\mathcal{H}_{3}$, and $\mathcal{H}_{4}$ in Table 1. In particular, we analyze the hierarchical triple merger pair GW190514-GW190521 with ($\mathcal{H}_{1}$) and without ($\mathcal{H}_{2}$) the EM counterpart ZTF19abanrhr. Because this candidate itself is not a high confidence counterpart [4, 5, 6, 7, 8]. Table 1 presents the Bayes factors and odds for various association combinations of GW190514, GW190521, and ZTF19abanrhr, indicating that a common origin shared by the three transients is strongly favored based on the observed data sets, in contrast to the associations involving only two of them. In particular, we find that compared with random coincidence ($\mathcal{H}_{0}$), the odds of GW190514 and GW190521 as a hierarchical triple merger without EM counterpart ($\mathcal{H}_{2}$) is $0.4:1$ (while that with EM counterpart ($\mathcal{H}_{1}$) is $199:1$). This suggests that the EM counterpart provides critical discriminatory power, lifting the likelihood of the hierarchical merger hypothesis by more than two orders of magnitude over the null hypothesis, implying that the EM counterpart is pivotal to breaking degeneracies between a true astrophysical signal and chance alignment. The odds of $\mathcal{H}_{1}$ is $\sim$14 ($\sim$117) times greater than the odds of $\mathcal{H}_{2}$, $\mathcal{H}_{3}$, and $\mathcal{H}_{4}$ (considering the primary masses). The analysis results, which account for the localizations and masses of the three transients, provide significantly robust support for the hypothesis of a common origin ($\mathcal{H}_{1}$). Moreover, we reanalyze the potential association by including the spins of the GW events. However, we find that the spins do not contribute significant information, yielding a Bayes factor of approximately 1.03 for the spin-only model.
@@ IV Conclusions and discussion
In this work, we investigate the potential association between GW190514 and GW190521 as a hierarchical triple merger associated with ZTF19abanrhr, taking into account of sky position, distance, and mass of the sources using a Bayesian criterion. Although the previous study has shown that GW190514 and GW190521 are a plausible hierarchical triple merger, their analysis is not significant with $p$-value of $\sim$0.14 [3] (generally, $p$-value of $\leq$0.05 is considered ‘significant’). Here, we introduce a new approach and propose using multi-messenger [4] signals to analyze the probability of the GW190514-GW190521 hierarchical triple merger. Our analysis (considering the primary masses) results greatly increase the probability, finding that a log Bayes factor of ${\rm log}\mathcal{B}=16.8~{}(21.5)$, corresponding to an odds ratio of $\mathcal{O}\sim$$1199~{}(2\times 10^{4}):$ 1, assuming an astrophysical prior odds of $\mathcal{P}\sim 10^{-5}$. This suggest strong evidence for the first hierarchical triple merger associated with an EM counterpart in the AGN formation channel. The finding of the multi-messenger hierarchical triple merger in the AGN is crucial for understanding the formation mechanisms of massive BHs, the role of AGNs in hierarchical mergers, and the implications of multi-messenger astronomy
@@
We show that the GW190514-GW190521 pair, which is separated by only one week, further supports the scenario of a triple hierarchical merger chain if associated with ZTF19abanrhr: i) Such a short time frame between the two mergers presents significant challenges for the reformation of a binary system and its subsequent merger. However, AGNs provide an environment conducive to overcoming these challenges [56, 57]. Specifically, the number density of BHs in the migration traps [36] of AGN disks is notably high due to migration [58]. This high density facilitates the rapid reformation of a binary system from GW190514’s remnant BHs, enabling a swift second merger with the aid of the dense gas in AGN disks; ii) GW190521 may not have had sufficient time to circularize in such a short interval, suggesting that it is likely to be eccentric [59, 60]. This inference is supported by previous analyses indicating a non-zero eccentricity for GW190521 [3], and AGN disk environments are known to promote an excess of eccentric mergers [17]. Our current methodological framework explicitly excludes time delay and eccentricity considerations in the statistical analysis, as these require dedicated waveform development that will be addressed in subsequent studies. Furthermore, we anticipate environmental effects to induce luminosity distance corrections of $\lesssim 1\%$—a systematic uncertainty. This is tiny compared to the uncertainty in intrinsic luminosity distance for each BBH observation (typically ±$\pm 30\%$). The impact warrants future investigation through targeted parameter estimation studies [61].
Our Bayesian analysis conditions on the association with ZTF19abanrhr, while we observe that multiple candidate EM counterparts to BBH mergers have been reported in the literature [21] since the first one for GW190521. This may introduce a potential trials factor that could affect the final significance of single association. However, our analysis focuses on evaluating the hierarchical merger scenario under the assumption that ZTF19abanrhr is a genuine EM counterpart. Incorporating a full population-level analysis to account for multiple GW-EM candidates is beyond the scope of this work but would be a valuable extension for future studies. The full analysis could include additional astrophysical properties of the AGN (beyond the three-dimensional spatial position of EM counterpart), such as the structure of the gas disk, the mass of the flare, etc, which may further confidently determine the validity of this association. Nevertheless, the high Bayes factor and odds ratio obtained here provide strong evidence for the plausibility of this particular hierarchical triple merger scenario, highlighting the importance of joint GW-EM analyses in probing BH formation channels.
@@ V Acknowledgments
We acknowledge valuable input from our anonymous referee which improved the original manuscript. This work is supported by National Key R&\&&D Program of China (2020YFC2201400), the National Natural Science Foundation of China (grant No. 11922303), and the Fundamental Research Funds for the Central Universities, Wuhan University (grant No. 2042022kf1182). We use the GW datasets released with GWTC-2.1 [2], labeled as “cosm” and “IMRPhenomXPHM.” This research has made use of data or software obtained from the Gravitational Wave Open Science Center (https://gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan. This analysis was made possible following software packages: NumPy [62], SciPy [63], Matplotlib [64], emcee [65], IPython [66], corner [67], seaborn [68], and Astropy [69].